OpenStudy (anonymous):

find two numbers with a sum of 20 and a difference of 14.

OpenStudy (anonymous):

x+y=20 x-y+14

OpenStudy (anonymous):

So sorry, x+y=20 x-y=14

OpenStudy (anonymous):

i meant to say the actual numbers that have a sum of 20 and a difference of 14

OpenStudy (anonymous):

Yes, the first number I put as x, and the other number I put as y so I could solve